Two large containers A and B of the same size are filled with different fluids. The fluids in containers A and B are maintained at 0° C and 100° C, respectively. A small metal bar, whose initial temperature is 100° C, is lowered into container A. After 1 minute the temperature of the bar is 90° C. After 2 minutes (since being lowered into container A) the bar is removed and instantly transferred into the other container. After 1 minute in container B the temperature of the bar rises 10°. How long, measured from the start of the entire process, will it take the bar to reach 99.4° C? (Round your answer to two decimal places. Assume the final temperature being asked for is reached while the bar is container B.)

t = ___ min
1
Expert's answer
2020-06-22T18:16:52-0400

The DE for Newton's Law of cooling /warming is dT/dt=k(T -"T_{m})" --> dT/"T- \nT_{m}" =kdt

"\therefore \int \frac{dT}{T-T_{m}}=\int kdt"


="ln\left | T-T_{m} \right |=kt+c_{1}"


"T-T_{m}=e^{kt+c_{1}}" =T="T_{m}+e^{c_{1}}e^{kt}" ="T_{m}+Ce^{kt}"

For the first process:


Since the surrounding medium is container A->"T_{m}=0^{0}C"


Now applying the initial condition T(0)= "100^{0}C" we get 100=0+C=>C=100=> T= "100e^{kt}"


since the temperature decreased to "90^{0}" C after one minute,

90=100"e^{k1}" => "e^{k1}" =90/100 => k1=ln(9/10)=-0.10536=>T=100"e^{-0.10536t}"


thus after 2 minutes we have T=100"e^{(-0.10536*2)}" ="81^{0}C"


For hte second process:

Since the surrounding medium is container B=> "T_{m}=\n100^{0}C" .

now we have initial temperature ="81^{0}C"

thus applying the initial condition we get 81=100+C => C= 81-100 =>C=-19.

therefore T=100-19"e^{kt}"

since the temperature increased by "10^{0}" C after one minute

81+10=100-19"e^{k2}"

"19e^{k2}=9 =>e^{k2}=9\/19 =>k2=ln(9\/19)=-0.7472"


Hence T=100-"19e^{-0.7472t}"


thus for the bar to reach 99.4 we have


99.4=100-"19e^{-0.7472t} =>e^{-0.7472t}=0.6\/19"

therefore -0.7472 t=ln (0.6/19)


Hence t= 4.62 minutes


thus the total time from the entire process to reach 99.4 is t= t1+t2=2+4.62=6.62 minutes.